40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun May 16 12:51:08 2010, kiwamu, Update, Green Locking, reflected beam at PD 
    Reply  Sun May 16 19:25:45 2010, Koji, Update, Green Locking, reflected beam at PD 
       Reply  Thu May 20 13:29:40 2010, kiwamu, Update, Green Locking, mode profile at PPKTP crystal PPKTPmode.png
          Reply  Thu May 20 14:18:59 2010, kiwamu, Update, Green Locking, mode profile of 40m cavity MMT40mcavity.png
             Reply  Wed May 26 04:14:21 2010, kiwamu, Update, Green Locking, locked OLTF_endPDH.png
                Reply  Wed May 26 14:38:02 2010, Koji, Update, Green Locking, locked 
                   Reply  Thu May 27 02:22:24 2010, kiwamu, Update, Green Locking, more details stimes.pngspectrum.pngETMX_small.png
                      Reply  Thu May 27 09:43:50 2010, rana, Update, Green Locking, more details 
                         Reply  Mon Jul 12 13:16:53 2010, kiwamu, Update, Green Locking, PZT feedback at X end arm_day_night.png
                      Reply  Wed Jun 9 11:13:11 2010, kiwamu, Update, Green Locking, lock with PDH box 
                         Reply  Fri Jun 11 11:54:42 2010, kiwamu, Update, Green Locking, end PDH with thermal feedback LPF.pngthermal_feedback.png
                            Reply  Fri Jun 11 13:32:28 2010, Koji, Update, Green Locking, end PDH with thermal feedback 
                         Reply  Sun Jun 20 18:28:25 2010, kiwamu, Update, Green Locking, Re: lock with PDH box 
                Reply  Wed Jul 21 19:11:18 2010, Dmass, Update, Green Locking, locked 
Message ID: 3059     Entry time: Wed Jun 9 11:13:11 2010     In reply to: 2997     Reply to this: 3065   3092
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: lock with PDH box 

A progress on the end PDH locking :

by using a modified PDH box the green laser on the X-end station is locked to the arm cavity.

So far the end PDH locking had been achieved by using SR560s, but it was not sophisticated filter.

To have a sophisticated filter and make the control loop more stable, the PDH box labeled "#G1" was installed instead of the SR560s.

After the installation the loop looks more stable than the before.

Some details about the modification of the PDH box will be posted later.

 

Although, sometimes the loop was unlocked because the sum-amp (still SR560) which mixes the modulation and the feedback signal going to the NPRO PZT was saturated sometimes.

Thus as we expected a temperature control for the laser crystal is definitely needed in order to reduce such big low frequency drive signal to the PZT.

ELOG V3.1.3-